Question:A calorimeter contains 75g of water at 15 degree celsius.when 50g water of 100 degree celcius is poured in the calorimeter ,the temperature of the mixture becomes 25degree C.calculate the water equivalent of the calorimeter ?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
(m + 0.075)*1 *10 = (m+0.050)*1*75
10m + 0.750 = 75m + 3.750
65m = 3
m = 3/65 kg
 
Regards
Arun (askIITians forum expert)
m =